### v2

# Risk-Averse Learning with Delayed Feedback

Abstract
In real-world scenarios, risk-averse learning is valuable for mitigating potential adverse outcomes. However, the delayed feedback makes it challenging to assess and manage risk effectively. In this paper, we investigate risk-averse learning using Conditional Value at Risk (CVaR) as risk measure, while incorporating feedback with random but bounded delays. We develop two risk-averse learning algorithms that rely on one-point and two-point zeroth-order optimization approaches, respectively. The dynamic regrets of the algorithms are analyzed in terms of the cumulative delay and the number of total samplings. In the absence of delay, the regret bounds match the established bounds of zeroth-order stochastic gradient methods for risk-averse learning. Furthermore, the two-point risk-averse learning outperforms the one-point algorithm by achieving a smaller regret bound. We provide numerical experiments on a dynamic pricing problem to demonstrate the performance of the algorithms.
{IEEEkeywords}
Online convex optimization, risk-averse, dynamic regret, random delay
1 Introduction
In online learning, the decision-maker sequentially updates its decision in a changing environment relying on historical information such as observations of previous actions and costs. However, in real-world scenarios, these observations may arrive with delays and impact system performance. The performance of online learning is typically measured through regret, which quantifies the performance gap between the algorithm’s decisions and a chosen benchmark. A widely used benchmark is static regret [1], which compares against the best fixed decision in hindsight. In contrast, dynamic regret compares the algorithm’s performance against a time-varying sequence of optimal decisions, providing a more relevant measure in non-stationary or evolving settings [2, 3]. Common metrics used to analyze dynamic regret include function variation [2] and the path length of the optimal decisions [3].
Unlike traditional online learning that focuses on optimizing expected outcomes, risk-averse online learning optimizes the tail of the loss distribution to mitigate catastrophic outcomes under uncertainties [4, 5]. Its application includes financial investment [6], power grid management [7], and robotics [8]. Among various risk measures, CVaR is widely adopted, as it offers a coherent risk assessment by considering the tail end of the loss distribution [9]. This makes CVaR particularly useful in high-stakes applications such as finance [10], healthcare [11], and autonomous systems [12]. Additionally, risk-averse learning often assumes access only to function evaluations, making zeroth-order optimization a natural choice for estimating the CVaR gradient. The pioneering work [1] applies zeroth-order algorithms to online convex optimization. This method evaluates the function at a single point per iteration, a procedure known as one-point zeroth-order optimization. Based on this, various approaches have been developed to enhance the estimation performance, including the one-point residual [13] and multi-point zeroth-order optimization [14, 15]. Specifically, the one-point residual zeroth-order optimization algorithm queries function value once at each iteration and estimates the gradient using the residual between two consecutive points [13]. The multi-point zeroth-order algorithm queries function values under multiple perturbed actions at each iteration and estimates the gradient based on differences between function values queried simultaneously [14, 15].
Many real-world learning systems operate in environments with delayed feedback. For instance, in online learning platforms, the impact of an educational intervention might be observed with a time lag. In medical diagnosis, the efficacy of a treatment plan might become apparent after several weeks or months. In online recommender systems, click-through rates are aggregated and reported back at regular intervals. Accounting for delays is essential, as relying on outdated information degrades system performance, especially in risk-averse settings where timely data is essential for making conservative decisions. Efforts have been devoted to online learning with different delay settings [16, 17, 18, 19, 20, 21, 22, 23, 24]. To name a few, [16] investigates online learning with feedback delayed by a constant number of steps and analyzes the regret bound of the algorithm concerning the constant delay. The paper [17] investigates bandit convex optimization and multi-armed bandit with delayed feedback and analyzes the regret bound of the algorithm concerning the learning horizon and the cumulative feedback delay. The paper [18] investigates bandit online learning in multi-agent games and develops no-regret algorithms that ensure convergence to Nash equilibrium when individual delays are tame. Some work, e.g., [20, 21], investigate multi-armed bandit learning with delayed feedback, where the delays are independent and identically distributed (i.i.d.) random variables with a finite expectation [20] and subexponential [21], respectively. Notably, the above works evaluate algorithm performance using static regret. However, since delayed feedback presents greater challenges in changing environments, the study of online learning with delays in dynamic settings is more relevant. In such contexts, dynamic regret serves as a more appropriate metric. Only a few works, such as [23, 24, 22], have studied online learning under the dynamic regret framework, addressing the fixed constant delay [23, 24] and arbitrary delay [22].
In this study, we investigate risk-averse learning with delayed feedback, where the delay is varying yet upper bounded. We assume access only to function evaluations and use arrival feedback to construct the empirical distribution of the stochastic costs. Based on this, we estimate the CVaR gradient using one-point and two-point zeroth-order optimization methods. The estimation error is bounded by the difference between the true and the constructed empirical distribution functions. Unlike existing literature [17], which assumes that all feedback arrives before the iteration horizon ends, we address the scenario where the delay may cause some feedback packets to be unavailable by the end of the horizon. Additionally, random delays might change the order of feedback sequence and introduce drift in the path length of the comparators. To address this issue, [22] assumes that delays are bounded and do not alter the order of feedback. In this work, we remove this In-Order assumption and provide a general regret bound under the arbitrary delay setting. To quantify the error introduced by delayed feedback in gradient descent, we sort the received feedback by their utilization order and align it along a virtual timeline. Then, we compute the cumulative delay, which is the aggregate of the delays experienced by the feedback. The dynamic regret bounds of the algorithms are analyzed in terms of the delay, the total sampling number, and the path length of comparators. By appropriately selecting the parameters, we theoretically show that: 1) the regret bound of the risk-averse learning algorithms decreases as the total sampling number increases or the delays decrease; 2) the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a lower regret bound.
The remainder of this article is structured as follows: Section 2 introduces preliminaries and problem formulation. Section 3 presents the main results on the one-point and two-point risk-averse learning with delayed feedback. Section 4 demonstrates the efficacy of the algorithm through numerical simulations. Section 5 draws conclusions.
Notations: Let $\|\cdot\|$ denote the $l_{2}$ norm. Let the notation $\mathcal{O}$ hide the constant and $\tilde{\mathcal{O}}$ hide constant and polylogarithmic factors of the number of iterations $T$, respectively. Let $A\oplus B=\{a+b|a\in A,b\in B\}$ denote the Minkowski sum of two sets of position vectors $A$ and $B$ in Euclidean space.
2 Preliminaries and problem statements
Consider the cost function $J_{t}(x,\xi):\mathcal{X}\times\Xi\rightarrow\mathbb{R}$ for $t=1,\dots,T$, where $x\in\mathcal{X}$ denotes the decision variable with $\mathcal{X}\subseteq\mathbb{R}^{d}$ being the admissible set, $\xi\subseteq\Xi$ denotes the random noise and $T$ denotes the horizon. Assume that $\mathcal{X}$ contains the ball of radius $r$ centered at the origin, i.e., $r\mathbb{B}\subseteq\mathbb{R}^{d}$. Denote the diameter of the admissible set $\mathcal{X}$ as $D_{x}=\sup_{x,y\in\mathcal{X}}\|x-y\|$.
2.1 CVaR
We use CVaR as the risk measure. Suppose $J_{t}(x,\xi)$ has the cumulative distribution function $F_{t}(y)=P(J_{t}(x,\xi)\leq y)$, and is bounded by $U>0$, i.e., $|J_{t}(x,\xi)|\leq U$. Given a confidence level $\alpha\in(0,1]$, the VaR value is defined as
$J_{t}^{\alpha}=F_{t}^{-1}(\alpha):=\inf\{y:F_{t}(y)\geq\alpha\}.$
The $\mathrm{CVaR}_{\alpha}$ measures the expected $\alpha$-fractional shortfall of $J_{t}(x,\xi)$, which is
$C_{t}(x)$ $:=\mathrm{CVaR}_{\alpha}\left[J_{t}(x,\xi)\right]$
$=\mathbb{E}_{F}\left[J_{t}(x,\xi)|J_{t}(x,\xi)\geq J_{t}^{\alpha}\right].$ (1)
Since the exact CVaR gradient for learning strategies is generally unavailable, we construct its estimate using a zeroth-order optimization approach that queries the objective function at perturbed actions [1]. We perturb the action $\hat{x}=x+\delta u$ and query the loss $J_{t}(\hat{x},\xi)$, where $u$ is a vector sampled from a unit sphere $\mathbb{S}^{d}\in\mathbb{R}^{d}$ and $\delta$ is the perturbation radius, also known as the smoothing parameter.
2.2 Delay model
Assume that the loss $J_{t}(\hat{x}_{t},\xi_{t})$ will be observed after $d_{t}$ time slots, where $d_{t}\geq 0$ can vary from slot to slot. Denote $J_{s|t}(\hat{x}_{s|t})$ as the loss incurred at time slot $s$ under the action $\hat{x}_{s}$ and be observed at time slot $t$. At each time slot, the learner collects the losses and the perturb direction vectors, which is denoted as $\mathcal{L}_{t}=\{J_{s|t}(\hat{x}_{s|t}),u_{s|t}|s+d_{s}=t,s\leq t\}$. When the learner does not receive loss at time $t$, we have $\mathcal{L}_{t}=\emptyset$. When the learner receives multiple losses at time $t$, we denote $\Delta_{t}$ as the number of arrival feedback at time $t$. Furthermore, we sort the received losses according to their generating time and denote them as $J_{s_{1}|t}(\hat{x}_{s_{1}|t}),\dots,J_{s_{\Delta_{t}}|t}(\hat{x}_{s_{\Delta_{t}}|t})$, where $s_{n}$, for $n=1,\dots,\Delta_{t}$, is the generating time of the arrival feedback. Then, the collection of the learner at time $t$ is $\mathcal{L}_{t}=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,s_{n}\leq t,n=1,\dots,\Delta_{t}\}$, see Fig. 1 for an illustrative example.
[…]
Figure 1: An example of delay model: the learner does not receive feedback at slot t1t-1, i.e., $\mathcal{L}_{t-1}=\emptyset$, and receives two feedbacks at slot $t$, which are generated at slots t2t-2 and $t$, respectively, i.e., $\mathcal{L}_{t}=\{J_{t-2|t},u_{t-2|t},J_{t|t},u_{t|t}\}$.
2.3 Problem statement
We consider the class of games that satisfy the following assumptions.
Assumption 1
The delay is bounded, i.e., $d_{t}\leq\bar{d}$, for $t\in\{1,\dots,T\}$, with $\bar{d}\geq 0$.
Assumption 2
The cost function $J_{t}(x,\xi)$ is convex in $x$ for every $\xi\in\Xi$.
Assumption 3
The cost function $J_{t}(x,\xi)$ is $L_{0}$-Lipschitz continuous in $x$ for every $\xi\in\Xi$, i.e., for all $x,y\in\mathcal{X}$, we have $|J_{t}(x,\xi)-J_{t}(y,\xi)|\leq L_{0}\|x-y\|$ with $L_{0}>0$.
We evaluate the performance of the algorithm through the dynamic regret, which is defined as the cumulative loss under the performed actions against best actions in hindsight:
$\mathrm{DR}(T)=\sum_{t=1}^{T}C_{t}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}),$ (2)
where $\hat{x}_{t}$ is the action generated by the algorithm at time slot $t$, and $x_{t}^{\ast}={\textrm{argmin}}_{x\in\mathcal{X}}C_{t}(x)$ is the optimal action at time $t$. This paper aims to design risk-averse learning algorithms that achieve a sublinear regret, i.e., $\lim_{T\rightarrow\infty}{\rm DR}(T)/T=0$.
3 Main result
Since only function evaluations are available, we use the zeroth-order optimization method to estimate the CVaR gradient for the learning process. A smoothed approximation [1] of the CVaR function is constructed as
$C_{t}^{\delta}(x)=\mathbb{E}_{u\sim\mathbb{S}^{d}}[C_{t}(x+\delta u)].$ (3)
Then, the gradient of the smoothed function (3) is given by
$\nabla C_{t}^{\delta}(x)=\mathbb{E}_{u\sim\mathbb{S}^{d}}[\frac{d}{\delta}f(x+\delta u)u].$ (4)
The following lemmas analyze the CVaR function and the smoothed function, which are prepared for the regret bound analysis.
Lemma 1
[4] Given Assumption 2, $C_{t}(x)$ and $C_{t}^{\delta}(x)$ are both convex in $x$.
Lemma 2
[4] Given Assumption 3, $C_{t}^{\delta}(x)$ is $L_{0}$-Lipschitz continuous in $x$, i.e., $|C_{t}^{\delta}(x)-C_{t}(x)|\leq\delta L_{0}$.
Lemma 3
[25] Let $F$ and $G$ be two cumulative distribution functions of random variables, where the random variables are bounded by $U$. Then, we have that
$|{\rm CVaR}_{\alpha}[F]-{\rm CVaR}_{\alpha}[G]|\leq\frac{U}{\alpha}\sup_{x}\big{|}F(x)-G(x)\big{|}.$ (5)
3.1 One-point risk-averse learning
Given the difficulty of estimating the CVaR gradient with only one single sample, we sample multiple stochastic costs to perform the estimation at each iteration. Denote $m_{t}$ as the sampling number at time slot $t$. Let the sampling number $m_{t}$ satisfy
$\sum_{t=1}^{T}\frac{1}{\sqrt{m_{t}}}\leq cT^{1-\frac{a}{2}},$ (6)
where $a,c>0$ are tuning parameters. An example of the sampling policy that satisfies (6) is provided in [25], which is given as
$m_{t}=\phi(t)=\left\lceil b(\Delta_{T}-t+1)^{a}\right\rceil,$ (7)
where $b>0$ is the tuning parameter. In (6) and (7), the total sampling number over the iteration horizon $T$ increases with $a$. At each time slot, we query the stochastic cost for $m_{t}$ times under the perturbed action and obtain $J(\hat{x}_{t},\xi^{i}),~i=1,\ldots,m_{t}$. We assume that the samples queried within the same time slot are collected as one batch. At time slot $t$, the learner receives
$\mathcal{L}_{t}$ $=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t},\xi_{s_{n}|t}^{i}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,$
$\hskip 20.00003pts_{n}\leq t,i=1,\ldots,m_{s_{n}},n=1,\dots,\Delta_{t}\}.$ (8)
Then, the empirical distribution function is constructed as:
$\hat{F}_{s_{n}|t}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{i=1}^{m_{s_{n}}}\mathbf{1}\{J(\hat{x}_{s_{n}},\xi_{s_{n}}^{i})\leq y\},$ (9)
for $n=1,\dots,\Delta_{t}$. Accordingly, by the CVaR definition (2.1), the CVaR estimate is denoted as ${\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}]$ and the corresponding CVaR gradient estimate is
$\hat{g}_{t}^{n}=\frac{d}{\delta}{\rm CVaR}_{\alpha}\big{[}\hat{F}_{s_{n}|t}\big{]}u_{s_{n}|t},$ (10)
for $n=1,\dots,\Delta_{t}$ and $t=1,\dots,T$. When the learner receives multiple feedbacks, the gradient descent processes as
$\begin{array}[]{rl}x_{t}^{n}&=\mathcal{P}_{\mathcal{X}^{\delta}}(x_{t}^{n-1}-\eta\hat{g}_{t}^{n}),\ {\rm for}~n=1,\dots,\Delta_{t},\\ x_{t+1}&=x_{t}^{\Delta_{t}},\end{array}$ (11)
where the initial value is $x_{t}^{0}=x_{t}$ and $\mathcal{P}_{\mathcal{X}^{\delta}}(x):={\textrm{argmin}}_{y\in\mathcal{X}^{\delta}}\|x-y\|^{2}$ denotes the projection operation with $\mathcal{X}^{\delta}=\{x\in\mathcal{X}|\frac{1}{1-\delta/r}x\in\mathcal{X}\}$ being the projection set. When the learner does not receive any feedback at time slot $t$, the risk-averse learning algorithm updates as $x_{k+1}=x_{k}$. The projection keeps the actions $\hat{x}_{t}$ within the admissible set $\mathcal{X}$ as $\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\delta\mathbb{B}=\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}r\mathbb{B}\subseteq\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}\mathcal{X}=\mathcal{X}$. Since the learner perturbs the action once at each time slot, we call it one-point risk-averse learning, which is summarized in Algorithm 1.
Algorithm 1 One-point risk-averse learning with delayed feedback
0: Initial value $x_{0}$, iteration horizon $T$, smoothing parameter $\delta$, risk level $\alpha$, step size $\eta$
1:for ${\rm{slot}}\;t=1,\dots,T$ do
2:  Select sampling number $m_{t}=\phi(t)$
3:  Sample $u_{t}\in\mathbb{S}^{d}$
4:  Play $\hat{x}_{t}=x_{t}+\delta u_{t}$
5:  for $i=1,\ldots,m_{t}$ do
6:   Query $J_{t}(\hat{x}_{t},\xi_{t}^{i})$
7:  end for
8:  Collect feedback $\mathcal{L}_{t}$, as in (3.1)
9:  if $\mathcal{L}_{t}=\emptyset$ then
10:   Update: $x_{t+1}=x_{t}$
11:  else
12:   Set $x_{t}^{0}=x_{t}$
13:   for  $n=1,\dots,\Delta_{t}$  do
14:    Build empirical distribution function $\hat{F}_{s_{n}|t}(y)$, as in (9)
15:    Estimate CVaR: ${\rm{CVaR}}_{\alpha}[\hat{F}_{s_{n}|t}]$
16:    Estimate CVaR gradient: $\hat{g}_{t}^{n}$, as in (10)
17:    Set $x_{t}^{n}=\mathcal{P}_{\mathcal{X}^{\delta}}(x_{t}^{n-1}-\eta\hat{g}_{t}^{n})$
18:   end for
19:   Update: $x_{t+1}=x_{t}^{\Delta_{t}}$
20:  end if
21:end for
To quantify the error introduced by using outdated feedback in gradient descent, it is crucial to specify the feedback utilized at each time slot. To do it, we map the received losses to virtual time slots according to their utilization order. Namely, the $\tau$-th virtual time slot is associated with the $\tau$-th batch of utilization feedback $\{\tilde{J}_{\tau}(\tilde{x}_{\tau},\tilde{\xi}_{\tau}^{i}),i=1,\dots,m_{\tau}\}$. We assume that delays may alter the arrival order of feedback. Correspondingly, as described in Section 2.2, we sort the feedback received in the same time slot by their generation times. This setting facilitates the regret analysis, as shown in Lemma 5 in the next section.
Note that $\Delta_{t}$ is the number of feedback received by the learner at time $t$, the overall number of feedback received by the learner at the end of time $t$ is $L_{t}=\sum_{k=1}^{t}\Delta_{k}$. Denoting $\tau:=L_{t-1}+1$, for some $\tau\in\{1,\dots,T\}$, we have $\tilde{J}_{\tau}(\cdot)=J_{s_{1}|t}(\cdot),\dots,\tilde{J}_{\tau+\Delta_{t}-1}(\cdot)=J_{s_{\Delta_{t}}|t}(\cdot)$. Fig. 2 provides an illustrative example of the mapping process. Moreover, we use $\mu(\tau)$ to map the $\tau$-th virtual time slot to the real time slot, i.e., $\tilde{J}_{\tau}(\cdot)=J_{\mu(\tau)|\mu(\tau)+d_{\mu(\tau)}}(\cdot)=J_{\mu(\tau)}(\cdot)$. In this study, we analyze the regret bound of the algorithm in terms of the cumulative delay and the iteration horizon. To compute the cumulative delay, we define an auxiliary variable $\tilde{s}_{\tau}=\tau-1-L_{\mu(\tau)-1}$.
In the following lemma, we calculate the cumulative delay, accounting for the possibility that some packets may not be received by the end of the horizon.
[…]
Figure 2: Mapping from the real timeline to the virtual timeline.
Lemma 4
The following claims hold: 1) $\tilde{s}_{\tau}\geq 0$, for $\tau=\{1,\dots,L_{T}\}$; 2) $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2}$.
The proof of Lemma 4 is provided in Appendix.
Remark 1
Some work, e.g., [17, 22], assume that all feedback arrives by the end of time slot $T$ or that the impact of losing final packets is negligible when the horizon $T$ is sufficiently large. This assumption yields $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{T}d_{t}$, which is the cumulative delay over time horizon $T$. According to Lemma 4, we define a new cumulative delay:
$D:=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2},$ (12)
which accounts for the practical scenario where some feedback may not arrive before time $T$.
Before presenting the main results, we define the path length of optimal decisions, also known as the path length of comparators, which serves as a measure for dynamic regret analysis.
Definition 1
The path length of the comparison sequence is defined as $\mathcal{P}_{T}=\sum_{t=1}^{T-1}\|x_{t}^{\ast}-x_{t+1}^{\ast}\|$, where $x_{t}^{\ast}=\arg\min_{x\in\mathcal{X}^{\delta}}C_{t}(x)$, for all $t$, is the optimal decision at time slot $t$.
Given Definition 1, the following theorem analyzes the regret bound of Algorithm 1.
Theorem 1
Let Assumptions 1, 2 and 3 hold. Let the sampling numbers over the iteration horizon $T$ satisfy (6). We have the following claims:
  1. 1.
    when $a<1$, selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3a}{4}}\sqrt{\bar{d}P_{T}}$. Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}\big{)}\sqrt{\bar{d}P_{T}}+(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D\Big{)}$ with high probability.
  2. 2.
    when $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ with high probability.
Proof. The regret (2) is written as
$\hskip-5.0pt\mathrm{R}_{1}(T)$ $=\sum_{t=1}^{T}\mathbb{E}_{u}[C_{t}(\hat{x}_{t}+\delta u-\delta u)]-\mathbb{E}_{u}[C_{t}(x_{t}^{\ast}+\delta u-\delta u)]$
$\leq\sum_{t=1}^{T}\Big{(}C_{t}^{\delta}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\ast})\Big{)}+2\delta L_{0}T$
$\leq\sum_{t=1}^{T}\Big{(}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\ast})\Big{)}+3\delta L_{0}T,$ (13)
where the first inequality is from the Lipschitzness of $C_{t}$ and the definition of $C_{t}^{\delta}$, as in (3). The second inequality is from the Lipschitzness of $C_{t}^{\delta}$. Furthermore, for $t=1,\dots,T$, we have
$\min_{x_{t}\in\mathcal{X}^{\delta}}C_{t}^{\delta}(x_{t})$ $=\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}((1-\delta/r)x_{t})$
$\leq\min_{x_{t}\in\mathcal{X}}(\delta/r)C_{t}^{\delta}(0)+(1-\delta/r)C_{t}^{\delta}(x_{t})$
$\leq\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}(x_{t})+(\delta/r)L_{0}\left\|x_{t}\right\|$
$\leq\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}(x_{t})+D_{x}L_{0}\delta/r,$ (14)
where the first inequality is from the convexity of $C_{t}^{\delta}$, as in Lemma 1, and the second inequality is from Lipschitzness of $C_{t}^{\delta}$, as in Lemma 2. Denote $x_{t}^{\delta,\ast}={\textrm{argmin}}_{x\in\mathcal{X}^{\delta}}C_{t}^{\delta}(x)$. Substituting (3.1) into (3.1), we have
$\mathrm{R}_{1}(T)\leq\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\delta,\ast})+(3+D_{x}/r)\delta L_{0}T,$ (15)
Denote $\tilde{C}_{\tau}(\cdot)=C_{\mu(\tau)}^{\delta}$. The first and second terms of (15) is further written as
$\hskip 10.00002pt\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\delta,\ast})$
$\leq\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)})-\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}^{\delta,\ast})+2U\bar{d}$
$\leq\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}+2U\bar{d}$
$=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-\tilde{x}_{\tau}+\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}$
$\hskip 10.00002pt+\tilde{x}_{\tau}^{\delta,\ast}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}+2U\bar{d},$
$=R_{1}+R_{2}+R_{3}+2U\bar{d}$ (16)
with $R_{1}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-\tilde{x}_{\tau}\big{\rangle}$, $R_{2}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\big{\rangle}$ and $R_{3}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),\tilde{x}_{\tau}^{\delta,\ast}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}$. In (3.1), the first inequality follows from $T-L_{T}\leq\bar{d}$ and $|C_{t}(\cdot)|\leq U$. The second inequality is from the convexity of $C_{t}$, for all $t$. In the following, we are going to bound $R_{1}$, $R_{2}$ and $R_{3}$.
Denote $\tilde{g}_{\tau}$ as the $\tau$-th CVaR gradient estimate being used for gradient descent (11), for $\tau=1,\dots,L_{T}$. Since $x_{t}=\tilde{x}_{L_{t-1}+1}$ and $x_{\mu(\tau)}=\tilde{x}_{L_{\mu(\tau)-1}+1}=\tilde{x}_{\tau-\tilde{s}_{\tau}}$, we have
$\hskip 10.00002pt\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|$
$=\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\|\tilde{x}_{\tau-\tilde{s}_{\tau}+j}-\tilde{x}_{\tau-\tilde{s}_{\tau}+j+1}\|$
$\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\eta\|\tilde{g}_{\tau-\tilde{s}_{\tau}+j}\|\leq\frac{d\eta}{\delta}U\tilde{s}_{\tau},$ (17)
where the second inequality is from $\|\tilde{x}_{i}-\tilde{x}_{i+1}\|=\|\tilde{x}_{i}-\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{i}-\eta\tilde{g}_{i})\|\leq\eta\|\tilde{g}_{i}\|$ and the last inequality is from $\|\tilde{g}_{i}\|\leq\frac{dU}{\delta}$, for $i\in\{\tau-\tilde{s}_{\tau},\dots,\tau-1\}$, according to its definition specified in (10). Thus, for $R_{1}$ in (3.1), we have
$R_{1}$ $\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}L_{0}\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\frac{\eta L_{0}dU}{\delta}\tilde{s}_{\tau}=\frac{\eta L_{0}dUD}{\delta},$ (18)
where the first inequality is Lemma 2, the second inequality is from substituting (3.1) into (3.1) and the equality is from the definition of cumulative delay as in (12).
The next is to bound $R_{2}$ in (3.1). The equality is from Lemma 4. By the gradient descent (11), we have
$\hskip 13.00005pt\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$=\|\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{\tau}-\eta\tilde{g}_{\tau})-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$\leq\|\tilde{x}_{\tau}-\eta\tilde{g}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$=\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}+\eta^{2}\|\tilde{g}_{\tau}\|^{2}-2\eta\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\rangle,$
where the inequality follows from the fact that $\tilde{x}_{\tau}^{\delta,\ast}\in\mathcal{X}^{\delta}$. Then, we obtain
$\hskip 10.00002pt\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\rangle$
$\leq\frac{\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}.$ (19)
Note that constructing the empirical distribution function of the cost function using finite samples induces CVaR estimate error $\hat{\epsilon}_{t}:={\mathrm{CVaR}}_{\alpha}[\hat{F}_{t}]-{\mathrm{CVaR}}_{\alpha}[F_{t}]$, where $\hat{F}_{t}$ is the empirical distribution function that corresponds to $F_{t}$, for $t=1,\dots,T$. From the gradient of smoothed function (4) and the definition of $\hat{g}_{t}^{n}$ given in (10), the error of $\tau$-th utilized CVaR gradient estimate is
$\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)})=\mathbb{E}[\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\epsilon}_{\tau}\tilde{u}_{\tau}].$ (20)
for $\tau=1,\dots,L_{T}$. Then, for $R_{2}$, we have
$R_{2}$ $=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\big{\langle}\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\frac{\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}\Big{]}$
$\hskip 10.00002pt+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|\Big{]}$
$\leq R_{21}+R_{22}+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta^{2}}L_{T},$ (21)
where $R_{21}=\frac{1}{{2\eta}}\sum_{\tau=1}^{L_{T}}\big{(}\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}\big{)}$ and $R_{22}=\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|]$. The first equality follows from substituting (20) into (3.1). The first inequality follows from substituting (3.1) into (3.1) and the second inequality follows from $\mathbb{E}[\|\tilde{g}_{\tau}\|^{2}]\leq\frac{d^{2}U^{2}}{\delta^{2}}$. Furthermore, we have
$R_{21}$ $=\frac{1}{2\eta}\sum_{\tau=1}^{L_{T}}\left(\|\tilde{x}_{\tau}\|^{2}-\|\tilde{x}_{\tau+1}\|^{2}+2(\tilde{x}_{\tau}^{\delta,\ast})^{\top}(\tilde{x}_{\tau+1}-\tilde{x}_{\tau})\right)$
$\leq\frac{1}{2\eta}\Big{(}D_{x}^{2}+2\sum_{t=2}^{L_{T}}\tilde{x}_{\tau}^{\top}(\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast})$
$\hskip 10.00002pt+2(\tilde{x}_{L_{T}}^{\delta,\ast})^{\top}\tilde{x}_{L_{T}+1}-2(\tilde{x}_{1}^{\delta,\ast})^{\top}\tilde{x}_{2}\Big{)}$
$\leq\frac{1}{2\eta}\left(3D_{x}^{2}+2D_{x}\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast}\|\right)$
$\leq\frac{1}{2\eta}\left(3D_{x}^{2}+2D_{x}\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|\right)$
$=\frac{1}{2\eta}\left(3D_{x}^{2}+4D_{x}\bar{d}P_{T}\right)$ (22)
where the first and the second inequalities are from the fact that the diameter of $\mathcal{X}$ is $D_{x}$. The third inequality is from the projection mapping and $\|\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast}\|=(1-\delta/r)\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ and the last equality is from Lemma 5.
The next is to bound $R_{22}$. Leveraging the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [26], we have
$\mathbb{P}\left\{\sup_{y}|\hat{F}_{t}(y)-F_{t}(y)|\geq\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}}\right\}\leq\bar{\gamma},$ (23)
with $m_{t}$ being the sampling number at time $t$. Denote the event in (23) as $A_{t}$, and denote $\mathbb{P}\{A_{t}\}$ as the occurrence probability of event $A_{t}$, for $t=1,\dots,L_{T}$. By Lemma 3, the estimation error (20) is bounded by
$|\hat{\varepsilon}_{t}|=\frac{U}{\alpha}\sup\big{|}\hat{F}_{t}-F_{t}\big{|}\leq\frac{U}{\alpha}\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}}$ (24)
with probability at least $1-\bar{\gamma}$, for $t=1,\dots,L_{T}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, by substituting (24) into $R_{22}$, we obtain that
$R_{22}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[|\tilde{\varepsilon}_{\tau}|]\leq\frac{dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2\tilde{m}_{\tau}}}$
$\leq\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (25)
with probability at least $1-\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{L_{T}}A_{\tau}\}\geq 1-\sum_{\tau=1}^{L_{T}}\mathbb{P}\{A_{\tau}\}=1-L_{T}\frac{\gamma}{L_{T}}=1-\gamma$ and $L_{T}\leq T$.
For $R_{3}$ in (3.1), we have
$R_{3}$ $=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\Big{\langle}\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\Big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\Big{(}\|\tilde{g}_{\tau}\|+\Big{\|}\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau}\Big{\|}\Big{)}\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$
$\leq R_{31}+R_{32}$ (26)
with $R_{31}=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\tilde{g}_{\tau}\|\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$ and $R_{32}=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau}\|\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$. The first equality follows from substituting (20) into (3.1). Similar to (3.1), we have
$R_{32}\leq\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}.$ (27)
Furthermore, for $R_{31}$, we have
$R_{31}$ $\leq\frac{dU}{\delta}\sum_{\tau=1}^{L_{T}}\|\tilde{x}_{\tau}^{\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\ast}\|\leq\frac{2dU}{\delta}\bar{d}P_{T}$ (28)
where the first inequality is from $\|\tilde{g}_{\tau}\|\leq\frac{dU}{\delta}$ and $\|\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\|=(1-\delta/r)\|\tilde{x}_{\tau}^{\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\ast}\|$. The second inequality is from Lemma 5.
Substituting (3.1), (3.1), (3.1), (3.1), (3.1), (27) and (28) into (15), with $L_{T}\leq T$, we have
${\rm R}_{1}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+\frac{\eta L_{0}dU}{\delta}D+\frac{\eta d^{2}U^{2}}{2\delta^{2}}T$
$\hskip-30.00005pt+\hskip-3.00003pt\frac{3D_{x}^{2}+4D_{x}\bar{d}P_{T}}{2\eta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2dU\bar{d}P_{T}}{\delta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (29)
with probability $1-\gamma$. When $a<1$, selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3a}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}\big{)}\sqrt{\bar{d}P_{T}}+(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D\Big{)}$ with probability $1-\gamma$. When $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ with probability $1-\gamma$. $\blacksquare$
Remark 2
In Theorem 1, it can be observed $T^{1-\frac{a}{4}}$, $(T+D)^{\frac{3a}{4}}$ and $(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D$ are all smaller than $(T+D)^{1-\frac{a}{4}}$ when $a<1$. Thus, we conclude that $\tilde{{\mathrm{R}}}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{1-\frac{a}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ is a more concise yet conservative compared to ${\rm DR}_{1}(T)$ in Theorem 1 when $a<1$. Additionally, it can be observed that the regret order of Algorithm 1 decreases as the sampling number parameter $a$ (i.e., the total number of samples over horizon $T$) increases and increases as the delay or path length of comparator increases.
Remark 3
The lower regret bound of one-point bandit learning in the delay-free case is $\tilde{\mathcal{O}}\Big{(}T^{\frac{3}{4}}\sqrt{1+P_{T}}\Big{)}$ [27]. In this work, we address the risk-averse learning and obtain a regret bound of $\tilde{\mathcal{O}}\big{(}(T+D)^{1-\frac{1}{4}\min{\{1,a\}}}\big{)}\sqrt{\bar{d}P_{T}}$. The dependence on the sampling parameter $a$ is from using queried cost values to construct the empirical distribution functions for the stochastic costs. When the number of samples is sufficiently large, i.e., $a$ is close to $1$, the term $(T+D)^{1-\frac{a}{4}}$ becomes $(T+D)^{\frac{3}{4}}$. In the absence of delay, the obtained regret bound then matches the result in [27]. Additionally, similar to [27], the dynamic regret bound exhibits a square-root dependence on the path-length, which becomes non-sublinear when $P_{T}\geq\sqrt{T}$. Nevertheless, in many practical cases, the path-length is relatively small.
3.2 Two-point risk-averse learning
In this section, we investigate the risk-averse learning using a two-point zeroth-order optimization method, which improves both the accuracy of CVaR estimation and the learning performance. At each time slot $t$, the stochastic costs are queried under two perturbed actions, i.e., $\hat{x}_{t}^{1}=x_{t}+\delta u_{t}$ and $\hat{x}_{t}^{2}=x_{t}-\delta u_{t}$, for $m_{t}$ times, which yields stochastic costs $J(\hat{x}_{t}^{1},\xi_{t}^{i})$ and $J(\hat{x}_{t}^{2},\xi_{t}^{j})$ for $i=1,\dots,m_{t}$ and $j=m_{t}+1,\dots,2m_{t}$. The sampling strategy $m_{t}=\phi(t)$ is the same as (6). Then, the performance of the algorithm is evaluated by the cumulative loss under the performed two actions against the best actions in hindsight, which is given by
$\mathrm{DR}_{2}(T)=\sum_{t=1}^{T}\frac{1}{2}\big{(}C_{t}(\hat{x}_{t}^{1})+C_{t}(\hat{x}_{t}^{2})\big{)}-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}).$ (30)
Assume that the samples generated simultaneously experience the same delay. At time slot $t$, the learner receives samples:
$\mathcal{L}_{t}$ $=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i}),J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{j}),u_{s_{n}|t}\ |$
$\hskip 20.00003pts_{n}+d_{s_{n}}=t,s_{n}\leq t,i=1,\ldots,m_{s_{n}},$
$\hskip 20.00003ptj=m_{s_{n}}+1,\dots,2m_{s_{n}},n=1,\dots,\Delta_{t}\}.$ (31)
The empirical distribution functions constructed by the collected delayed feedback are given by
$\hskip-10.00002pt\hat{F}_{s_{n}|t}^{1}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{i=1}^{m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i})\leq y\},$
$\hskip-10.00002pt\hat{F}_{s_{n}|t}^{2}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{j=m_{s_{n}}+1}^{2m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{j})\leq y\},$ (32)
for $n=1,\dots,\Delta_{t}$ and $t=1,\dots,T$. Accordingly, the CVaR estimates are ${\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}^{1}]$ and ${\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}^{2}]$, and the two-point CVaR gradient estimate is given as
$\bar{g}_{t}^{n}=\frac{d}{2\delta}\Big{(}{\rm CVaR}_{\alpha}\big{[}\hat{F}_{s_{n}|t}^{1}\big{]}-{\rm CVaR}_{\alpha}\big{[}\hat{F}_{s_{n}|t}^{2}\big{]}\Big{)}u_{s_{n}|t}.$ (33)
The learning process is the same as (11), where the gradient estimate is replaced with $\bar{g}_{t}^{n}$. The two-point risk-averse learning algorithm with delayed feedback is summarized in Algorithm 2.
Algorithm 2 Two-point risk-averse learning with delayed feedback
0: Initial value $x_{0}$, iteration horizon $T$, smoothing parameter $\delta$, risk level $\alpha$, step size $\eta$.
1:for ${\rm{slot}}\;t=1,\dots,T$ do
2:  Select sampling number $m_{t}=\phi(t)$
3:  Sample $u_{t}\in\mathbb{S}^{d}$
4:  Play $\hat{x}_{t}^{1}=x_{t}+\delta u_{t}$ and $\hat{x}_{t}^{2}=x_{t}-\delta u_{t}$
5:  for $i=1,\ldots,m_{t}$ and $j=m_{t}+1,\dots,2m_{t}$ do
6:   Query $J_{t}(\hat{x}_{t}^{1},\xi_{t}^{i})$ and $J_{t}(\hat{x}_{t}^{2},\xi_{t}^{j})$
7:  end for
8:  Collect feedback $\mathcal{L}_{t}$, as in (3.2)
9:  if $\mathcal{L}_{t}=\emptyset$ then
10:   Update: $x_{t+1}=x_{t}$
11:  else
12:   Set $x_{t}^{0}=x_{t}$
13:   for  $n=1,\dots,\Delta_{t}$  do
14:    Build empirical distribution functions $\hat{F}_{s_{n}|t}^{1}(y)$ and $\hat{F}_{s_{n}|t}^{2}(y)$, as in (3.2)
15:    Estimate ${\rm{CVaR}}_{\alpha}[\hat{F}_{s_{n}|t}^{1}]$ and ${\rm{CVaR}}_{\alpha}[\hat{F}_{s_{n}|t}^{2}]$
16:    Estimate CVaR gradient $\bar{g}_{t}^{n}$, as in (33)
17:    Set $x_{t}^{n}=\mathcal{P}_{\mathcal{X}^{\delta}}(x_{t}^{n-1}-\eta\bar{g}_{t}^{n})$
18:   end for
19:   Update: $x_{t+1}=x_{t}^{\Delta_{t}}$
20:  end if
21:end for
The following theorem analyzes the regret bound of Algorithm 2.
Theorem 2
Let Assumptions 1, 2 and 3 hold. Let the sampling numbers over the iteration horizon $T$ satisfy (6). Selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{1}{2}}\sqrt{\bar{d}P_{T}}$, Algorithm 2 achieves ${\rm DR}_{2}(T)=\tilde{\mathcal{O}}\bigg{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\big{)}\sqrt{\bar{d}P_{T}}\bigg{)}$ with high probability.
Proof. Firstly, we have
${\rm DR}_{2}(T)$ $\hskip-1.99997pt=\hskip-1.99997pt\sum_{t=1}^{T}\big{(}C_{t}(x_{t}+\delta u_{t})+C_{t}(x_{t}-\delta u_{t})\big{)}/2-\hskip-3.00003pt\sum_{t=1}^{T}C_{t}(x_{t}^{\ast})$
$=\sum_{t=1}^{T}C_{t}(x_{t}+\delta u_{t})-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}),$ (34)
where the first equality is from the definition of $\hat{x}^{1}$ and $\hat{x}^{2}$. The second equality is from the symmetric distribution of $u_{t}$. The proof follows the same approach as in Theorem 1. The derivation differs starting from (3.1). The CVaR gradient estimate (33) is bounded by
$\mathbb{E}[\|\bar{g}_{t}\|]$ $\leq\mathbb{E}_{u}\Big{[}\frac{d}{2\delta}\Big{\|}C_{t}(x_{t}+\delta u_{t})-C_{t}(x_{t}-\delta u_{t})\Big{\|}\big{\|}u_{t}\big{\|}\Big{]}$
$\leq L_{0}d,$ (35)
for all $t$, where the second inequality is from the Lipschitzness of $C_{t}$. Similar to (3.1), we have
$\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\eta\|\tilde{\bar{g}}_{\tau-\tilde{s}_{\tau}+j}\|\leq L_{0}d\eta\tilde{s}_{\tau}$ (36)
where $\tilde{\bar{g}}_{i}$ is $i$-th utilized two-point CVaR gradient estimate. Thus, for $R_{1}$ in (3.1), we have
$R_{1}$ $\leq\sum_{\tau=1}^{L_{T}}L_{0}\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\leq L_{0}^{2}d\eta\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=L_{0}^{2}d\eta D,$ (37)
where the first inequality is from the Lipschitzness of $C_{t}$. The second inequality is from substituting (36) into (37). The last inequality is from the definition of cumulative delay as in (12).
Additionally, in the two-point case, the error of the CVaR gradient estimate (33) is given as
$\bar{\varepsilon}_{t}$ $:={\rm CVaR}_{\alpha}[\hat{F}_{t}^{1}]-{\rm CVaR}_{\alpha}[\hat{F}_{t}^{2}]$
$\hskip 10.00002pt-\big{(}{\rm CVaR}_{\alpha}[F_{t}^{1}]-{\rm CVaR}_{\alpha}[F_{t}^{2}]\big{)},$ (38)
where $\hat{F}_{t}^{1}$ and $\hat{F}_{t}^{2}$ the empirical distribution functions at time $t$ and $F_{t}^{1}$ and $F_{t}^{2}$ denote the corresponding true distribution function. By the DKW inequality, we have that
$\mathbb{P}\left\{\sup_{y}|\hat{F}_{t}^{i}(y)-F_{t}^{i}(y)|\geq\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}}\right\}\leq\bar{\gamma},$ (39)
for $i=1,2$. Denote the event in (39) as $B_{t}^{i}$, and denote $\mathbb{P}\{B_{t}^{i}\}$ as the occurrence probability of event $B_{t}^{i}$, for $i=1,2$ and $t=1,\dots,L_{T}$. By Lemma 3, the CVaR estimate error (3.2) is bounded by
$|\bar{\varepsilon}_{t}|\leq\frac{2U}{\alpha}\sup\big{|}\hat{F}_{t}-F_{t}\big{|}\leq\frac{2U}{\alpha}\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}},$ (40)
with probability at least $1-\mathbb{P}\{\bigcup_{i=1}^{2}B_{t}^{i}\}=1-2\bar{\gamma}$, for $t=1,\dots,L_{T}$. Similar to (3.1), we have
$R_{2}$ $\leq R_{21}+\bar{R}_{22}+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\frac{\eta}{2}\|\tilde{\bar{g}}_{\tau}\|^{2}\Big{]}$
$\leq R_{21}+\bar{R}_{22}+\frac{\eta}{2}L_{0}^{2}d^{2}L_{T}.$ (41)
with $\bar{R}_{22}=\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\frac{d}{\delta}{\tilde{\bar{\varepsilon}}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|]$ and $\tilde{\bar{\varepsilon}}_{\tau}=\bar{\varepsilon}_{\mu(\tau)}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, we obtain
$\bar{R}_{22}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[|\tilde{\bar{\varepsilon}}_{\tau}|]\leq\frac{2dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2\tilde{m}_{\tau}}}$
$\leq\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (42)
with probability at least $1-2\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{T}\bigcup_{i=1}^{2}B_{\tau}^{i}\}\geq 1-\sum_{\tau=1}^{T}\sum_{i=1}^{2}\mathbb{P}\{B_{\tau}^{i}\}\geq 1-2\gamma$. The second inequality is from substituting (40) into (3.2).
Similar to (3.1)-(28), we have
$R_{3}$ $\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}(\|\tilde{\bar{g}}_{\tau}\|+\|\frac{d}{\delta}\tilde{\bar{\varepsilon}}_{\tau}\tilde{u}_{\tau}\|)\|\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\|\Big{]}$
$\leq 2\frac{dU}{\delta}\bar{d}P_{T}+\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}$ (43)
where the second inequality follows the derivation of (27) and (28). Substituting (3.1), (3.1), (37), (3.2), (3.2) and (3.2) into (3.2), and with $L_{T}\leq T$, we have
${\rm DR}_{2}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+L_{0}^{2}d\eta D+\frac{\eta}{2}L_{0}^{2}d^{2}T$
$\hskip-35.00005pt+\hskip-3.00003pt\frac{3D_{x}^{2}+4D_{x}\bar{d}P_{T}}{2\eta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2dU\bar{d}P_{T}}{\delta}\hskip-1.99997pt+\hskip-1.99997pt\frac{4cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}.$ (44)
Selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{1}{2}}\sqrt{\bar{d}P_{T}}$, Algorithm 2 achieves ${\rm DR}_{2}(T)=\tilde{\mathcal{O}}\bigg{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\big{)}\sqrt{\bar{d}P_{T}}\bigg{)}$ with probability $1-2\gamma$. $\blacksquare$
Remark 4
By comparing the regret bounds of Algorithms 1 and 2, i.e., $\mathrm{DR}_{1}(T)$ and $\mathrm{DR}_{2}(T)$ in Theorem 1 and 2, it can be observed that the two-point risk-averse learning algorithm (Algorithm 2) outperforms the one-point risk-averse learning algorithm (Algorithm 1), by achieving a smaller regret bound, for all $a>0$. Additionally, by comparing the order of the cumulative delay $D$, one finds that Algorithm 2 has a higher tolerance for delay compared to Algorithm 1.
4 Simulation
In this section, we consider the dynamic pricing problem for parking lots, see [28]. Factors such as parking prices, availability, and locations generally influence driving decisions. This encourages us to dynamically adjust the parking price according to real-time demand. Denote $r_{t}\in[0,1]$ as curb occupancy rate. The occupancy rate is influenced by the price $x_{t}$ and environmental uncertainties $\xi_{t}$: $r_{t}=\xi_{t}+Ax_{t}$, where A=0.15A=-0.15 is the estimated price elasticity, which is determined by [28] through analyzing the real-world data. Assume that the uncertainty $\xi$ follows a uniform distribution between $[0.9,1.1]$, i.e., $\mathcal{U}$. Generally, it is easier to find a parking space when the occupancy rate is maintained at $70\%$. Thus, we modeled the desired occupancy as $r_{d}=0.7$. To prevent either overcrowding or insufficient occupancy, our goal is to minimize the risk-averse objective function
[…]
where the risk level is selected as $\alpha=0.5$ and the loss function is given by
$J(x_{t},\xi_{t})=\|\xi_{t}+Ax_{t}-r_{d}\|^{2}+\frac{\nu}{2}\|x_{t}\|^{2},$
where $\nu=0.001$ is the regularization parameter. We set the initial price value as $x_{0}=1$ and restrict the potential prices to the range $[1,5]$.
Figure 3: From top to bottom: parking prices generated by the one-point algorithm (Algorithm 1) and by the two-point algorithm (Algorithm 2); the corresponding resulting occupancies and the desired occupancy rate $r_{d}=0.7$.
Figure 4: From top to bottom: the CVaR values under the price $\hat{x}_{t}$ generated by Algorithm 1 and under the prices $\hat{x}_{t}^{1}$ and $\hat{x}_{t}^{2}$ generated by Algorithm 2; the cumulated CVaR value achieved by the one-point algorithm (Algorithm 1) and the two-point algorithm (Algorithm 2).
Figure 5: From top to bottom: parking prices generated by the one-point algorithm (Algorithm 1) and by the two-point algorithm (Algorithm 2); the corresponding resulting occupancies and the desired occupancy rate $r_{d}=0.6$ when $t\in[0,200]$ and $r_{d}=0.7$ when $t\in(200,500]$.
Figure 6: From top to bottom: the CVaR values under the price $\hat{x}_{t}$ generated by Algorithm 1, under the prices $\hat{x}_{t}^{1}$ and $\hat{x}_{t}^{2}$ generated by Algorithm 2; the cumulated CVaR value achieved by the one-point algorithm (Algorithm 1) and the two-point algorithm (Algorithm 2).
We assume that the parking lot periodically monitors the occupancy rate and adjusts prices accordingly. Assume that the unit of time for an iteration is one week, which means the parking lot observes the occupancy rate and changes prices every week. However, due to data aggregation or computational overhead, these observations may be delayed. Assume that the random delays obey a uniform distribution with the range of $[0,5]$, with a maximum delay of $\bar{d}=5$. Namely, the maximum information delay is $5$ weeks. Set the sample number as $n_{t}=8$ for all $t$. Shaded areas represent $\pm$ one standard deviation over $20$ runs. Fig. 3 depicts parking prices and the corresponding occupancies generated by Algorithms 1 and 2. Fig. 3 demonstrates that the prices and occupancies generated by Algorithm 2 track the desired values faster compared to Algorithm 1. Fig. 4 depicts CVaR values and the dynamic regret $\mathrm{DR}(T)$ generated by Algorithms 1 and 2, respectively. Fig. 4 shows that the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving lower regret.
Additionally, we consider the scenario where the desired occupancy changes over time due to variations in demand or usage patterns. Assume that the desired occupancy rate is $r_{d}=0.65$ when $t\in[0,200]$ and $r_{d}=0.7$ when $t\in(200,500]$. Fig. 5 depicts parking prices and the corresponding occupancies generated by Algorithms 1 and 2, which demonstrates that the prices and occupancies generated by Algorithm 2 track desired values faster compared to Algorithm 1. Fig. 6 depicts CVaR values and the dynamic regret $\mathrm{DR}(T)$ of Algorithms 1 and 2 when facing the changing desired occupancy, respectively. Fig. 6 shows that the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a lower regret. Moreover, Figs. 3-6 reveal that Algorithm 2 yields results with a lower variance compared to Algorithm 1. This suggests that Algorithm 2 produces more consistent and stable outcomes than Algorithm 1.
Figure 7: The occupancies and cumulated losses generated by Algorithm 2 under maximal individual delays $\bar{d}=\{0,20,50\}$.
Figure 8: The occupancies and cumulated losses generated by Algorithm 2 with sample strategies $n_{t}=\{1,2,16\}$.
Fig. 7 depicts the resulting occupancies and the regret achieved by the two-point risk-averse learning algorithm under the maximum delay of $\bar{d}=\{0,20,50\}$. The sampling number is set as $n_{t}=8$. Fig. 7 shows that, the occupancy under Algorithm 2 converges to the desired value slower as the delay increases, and the variance of the resulting occupancy and the cumulated loss both increase. This effect is particularly pronounced when $\bar{d}=50$.
In Fig. 8, we fix the maximal delay as $\bar{d}=5$ and depict the resulting occupancies and the regret of Algorithm 2 with sampling number $n_{t}=\{1,2,16\}$ for all $t$. Fig. 8 shows that, as the sampling number increases, the occupancy under Algorithm 2 converges to the desired value faster and the variance of the resulting occupancy and the cumulated loss both decrease.
5 Conclusion
This paper investigated risk-averse learning with delayed feedback, where the delays are random yet bounded. We developed the one-point and two-point risk-averse learning algorithms, where the one-point and two-point zeroth-order optimization approaches are employed to estimate the CVaR gradient, respectively. In both cases, function values are queried multiple times under perturbed actions to improve the accuracy of the CVaR gradient estimate. To capture the error induced by using outdated feedback to process the gradient descent, we sorted the arriving feedback according to their arrival order and calculated the cumulative delay. The dynamic regret bounds of these algorithms are then analyzed in terms of the cumulative delay, the total sampling number and the path length of comparators. The theoretical results demonstrated that both the one-point and two-point risk-averse learning algorithms achieve sublinear regrets. Furthermore, the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a smaller regret bound.
6 Appendix
Proof of Lemma 4. The first claim is from [17], we provide its proof here for completeness. Denote $L_{\mu(\tau)-1}=m$, which means that the learner receives $m$ batches feedback before $t_{1}=\mu(\tau)$. Due to the delay, we have $0\leq m\leq\mu(\tau)-1$. The learner receives the $\tau$-th batch feedback at slot $t_{2}=\mu(\tau)+d_{\mu(\tau)}\geq t_{1}$. In other words, the learner receives at least $m$ batches feedback before $t_{2}$. Then, we have $\tau\geq m+1$, and further that $\tilde{s}_{\tau}\geq m+1-1-m=0$. Additionally, before the slot $\mu(\tau)$, the learner receives $L_{\mu(\tau)-1}$ batches feeback. Note that $L_{\mu(\tau)-1}\geq\mu(\tau)-\bar{d}-1$, we have $\tilde{s}_{\tau}=\tau-1-L_{\mu(\tau)-1}\leq\tau-1-t(\tau)+1+\bar{d}\leq 2\bar{d}.$
We next prove the second claim. According to the definition of $\tilde{s}_{\tau}$, we have
$\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}$ $=\sum_{\tau=1}^{L_{T}}\big{(}\tau-1-L_{\mu(\tau)-1}\big{)}$
$=\sum_{t=1}^{L_{T}}(t-1)-\sum_{\tau=1}^{L_{T}}L_{\mu(\tau)-1}$
$=\sum_{t=1}^{L_{T}}(t-1)-\sum_{t=1}^{L_{T}-\bar{d}}L_{t-1}-\sum_{\tau\in\mathcal{S}}L_{\mu(\tau)-1},$ (46)
with $\mathcal{S}=\{\tau|1\leq\tau\leq L_{T},\mu(\tau)>L_{T}-\bar{d}\}$. For the first and the second terms of (6), we have
$\hskip 10.00002pt\sum_{t=1}^{L_{T}}(t-1)-\sum_{t=1}^{L_{T}-\bar{d}}L_{t-1}$
$=\sum_{t=1}^{L_{T}-2\bar{d}-1}t-\sum_{t=1}^{L_{T}-\bar{d}-1}L_{t}+\sum_{t=L_{T}-2\bar{d}}^{L_{T}-1}t$
$\leq\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\sum_{t=L_{T}-2\bar{d}}^{L_{T}-1}t$
$=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}(2L_{T}-2\bar{d}-1),$ (47)
where the first equality is from $L_{0}=0$. We next explain the inequality. In the absence of delay, the learner receives $t$ batches of feedback by the end of slot $t$. Then, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}t$ can be interpreted as each packet is accounted once at each slot since its arrival over the horizon $L_{T}-\bar{d}-1$. Accordingly, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}L_{t}$ means that each packet is accounted once at each slot since its arrival in the presence of delay. Additionally, due to the delay, each $J_{t}(\cdot)$ is undercounted $d_{t}$ times by slot $t+\bar{d}$. By Assumption 1, the losses $J_{t}(\cdot)$, for $t\in\{1,\dots,L_{T}-2\bar{d}-1\}$, arrive before the end of $L_{T}-\bar{d}-1$ slot. Thus, we have $\sum_{t=1}^{L_{T}-2\bar{d}-1}t-\sum_{t=1}^{L_{T}-\bar{d}-1}L_{t}\leq\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}$. For the last term of (6), we have
$\sum_{\tau\in\mathcal{S}}L_{\mu(\tau)-1}$ $\geq\sum_{\tau\in\mathcal{S}}\mu(\tau)-\bar{d}-1\geq\sum_{L_{T}-2\bar{d}}^{L_{T}-\bar{d}-1}t$
$=\frac{\bar{d}(2L_{T}-3\bar{d}-1)}{2},$ (48)
where the first inequality is from $L_{t}\geq t-\bar{d}$ and the second inequality is from the fact the set $\mathcal{S}$ has $\bar{d}$ components and $\mu(\tau)\geq L_{T}-\bar{d}+1$ for $\tau\in\mathcal{S}$, as in (6). Substitute (6) and (6) into (6), we obtain the second claim. $\blacksquare$
Delays result in out-of-order feedback, which imposes challenges in comparing the decisions generated by the algorithm against the optimal ones. The following lemma is prepared for the regret analysis in Theorems 1 and 2.
Lemma 5
Consider the virtual mapping $\tilde{x}_{\tau}=x_{\mu(\tau)}$, the following inequalities hold.
$\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq 2\bar{d}P_{T}$ (49)
$\sum_{\tau=2}^{L_{T}}\|x_{\mu(\tau)}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq 2\bar{d}P_{T}$ (50)
Proof of Lemma 5. For some $\mu(\tau)=t$, we have $t-\bar{d}-1\leq\mu(\tau-1)\leq t+\bar{d}-1$. Then, for $t=\mu(\tau)>\bar{d}+1$, we have
$\hskip 10.00002pt\|\tilde{x}_{\tau-1}-\tilde{x}_{\tau}\|=\|x_{\mu(\tau-1)}-x_{\mu(\tau)}\|$
$\leq\max_{i\in\{t-\bar{d}-1,\dots,t+\bar{d}-1\}}\|x_{i}-x_{t}\|\leq\sum_{k=t-\bar{d}}^{t+\bar{d}-1}\|x_{k-1}-x_{k}\|.$
For $1\leq\mu(\tau)<\bar{d}+1$, we have $1\leq\mu(\tau-1)\leq t+\bar{d}-1$. Then, we have
$\|\tilde{x}_{\tau-1}-\tilde{x}_{\tau}\|\hskip-1.99997pt\leq\hskip-1.99997pt\max_{i\in\{1,\dots,t+\bar{d}-1\}}\|x_{i}-x_{t}\|\hskip-1.99997pt\leq\hskip-1.99997pt\sum_{k=2}^{t+\bar{d}-1}\|x_{k-1}-x_{k}\|,$
Furthermore, $\mu(\tau)\leq T$. Thus, we obtain
$\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq\sum_{t=2}^{T}\sum_{k=\min\{2,t-\bar{d}\}}^{t+\bar{d}-1}\|x_{k-1}^{\ast}-x_{k}^{\ast}\|\leq 2\bar{d}P_{T}$ (51)
The next is to prove (50). To analyze the range of $\mu(\tau)$, we consider the following two extreme cases. First, consider a real-time slot $t=k$. Under the assumption of a maximum delay of $\bar{d}$, at most $\bar{d}-1$ pieces of feedback have not been received. More specifically, suppose that the feedback $C_{k}$ is received by the learner without delay, while the feedbacks $\{C_{k-\bar{d}+1},\dots,C_{k-1}\}$ are all delayed by $\bar{d}$ time slots. In this case, for $\tau=k-\bar{d}+1$, we have $\mu(\tau)=k$. This yields $\mu(\tau)\geq\tau-\bar{d}+1$. Second, consider the opposite scenario where $C_{t}$ is delayed by $\bar{d}$ step, and thus received at time $t+\bar{d}$. By then, the learner has received at most $k+\bar{d}-1$ feedbacks. In this case, for $\tau=k+\bar{d}$, we have $\mu(\tau)=k$. This yields $\mu(\tau)\leq\tau+\bar{d}$. Combining both bounds, we obtain that $\mu(\tau)\in[\tau-\bar{d}+1,\tau+\bar{d}]$ and $x_{\mu(\tau)}\in\{x_{\tau-\bar{d}+1},\dots,x_{\tau+\bar{d}}\}$. Following similar arguments, one can prove $\tilde{x}_{\tau}\in\{x_{\tau-\bar{d}+1},\dots,x_{\tau+\bar{d}}\}$. Therefore, we have
$\hskip 10.00002pt\sum_{\tau=2}^{L_{T}}\|x_{\mu(\tau)}^{\ast}-\tilde{x}_{\tau}^{\ast}\|\leq\sum_{\tau=2}^{L_{T}}\max_{i,j\in\{\tau-\bar{d}+1,\tau+\bar{d}\}}\|x_{i}^{\ast}-x_{j}^{\ast}\|$
$\leq\sum_{\tau=2}^{L_{T}}\sum_{k=\min\{2,\tau-\bar{d}\}}^{\tau+\bar{d}}\|x_{k-1}^{\ast}-x_{k}^{\ast}\|\leq 2\bar{d}P_{T}$ (52)
where the last equality is from Definition 1. $\blacksquare$